# More Text, Less Point: Towards 3D Data-Efficient Point-Language Understanding

### v2

@@ Abstract
Enabling Large Language Models (LLMs) to comprehend the 3D physical world remains a significant challenge. Due to the lack of large-scale 3D-text pair datasets, the success of LLMs has yet to be replicated in 3D understanding. In this paper, we rethink this issue and propose a new task: 3D Data-Efficient Point-Language Understanding. The goal is to enable LLMs to achieve robust 3D object understanding with minimal 3D point cloud and text data pairs. To address this task, we introduce GreenPLM, which leverages more text data to compensate for the lack of 3D data. First, inspired by using CLIP to align images and text, we utilize a pre-trained point cloud-text encoder to map the 3D point cloud space to the text space. This mapping leaves us to seamlessly connect the text space with LLMs. Once the point-text-LLM connection is established, we further enhance text-LLM alignment by expanding the intermediate text space, thereby reducing the reliance on 3D point cloud data. Specifically, we generate 6M free-text descriptions of 3D objects, and design a three-stage training strategy to help LLMs better explore the intrinsic connections between different modalities. To achieve efficient modality alignment, we design a zero-parameter cross-attention module for token pooling. Extensive experimental results show that GreenPLM requires only 12% of the 3D training data used by existing state-of-the-art models to achieve superior 3D understanding. Remarkably, GreenPLM also achieves competitive performance using text-only data. The code and weights are available at: https://github.com/TangYuan96/GreenPLM.
footnotetext: * Equal contribution; $\dagger$ corresponding author.
@@ Introduction
To address this data-limited multimodal alignment problem, we propose GreenPLM. Intuitively, as shown in Fig. 2, we observe that after establishing the point-text-LLM connection, instead of increasing point-text data pairs to optimize the point-text mapping like in existing methods (Xu et al. 2023; Qi et al. 2024a), we can also enhance the text-LLM alignment by simply adding more text data. This approach can also improve the point-LLM alignment and, more importantly, reduce the reliance on point-text data pairs, shifting the data bottleneck from expensive and scarce 3D-text data to abundant and cheap text data. That is, the text-LLM alignment approach fits perfectly with the goal of 3D data-efficient point-language understanding, also provides an alternative solution for aligning point clouds with LLMs, enabling GreenPLM to achieve robust 3D understanding even with limited 3D data.
@@
Caption - A 3D model of a friendly white dog…
@@
The encoder maps raw data into a more compact embedding space, which can then be aligned with LLMs. To reduce the training cost, one can intuitively employ a multimodal pre-trained encoder, such as CLIP (Radford et al. 2021), which has been trained on text-image pairs, for aligning 2D images with LLMs. This makes it easier to align data from different modalities with LLMs. Similarly, some existing 3D-LLMs use multimodal pre-trained encoders (Huang et al. 2023; Xue et al. 2023; Qi et al. 2023; Gao et al. 2024; Lei et al. 2023; Chen et al. 2024a; Han et al. 2024) to map point clouds into embedding space, followed by fine-tuning the 3D-LLM. However, even without training the encoder, constructing the 3D-LLM still requires a vast amount of point-text data (Xu et al. 2023; Zhou et al. 2023; Qi et al. 2024a; Tang et al. 2024). We observe that existing methods underutilize the potential of the text encoder, only focusing on aligning point encoder with LLM. In contrast, we propose leveraging the cost-efficient text space and the text encoder to reduce the dependency on 3D data.
@@
Notably, to enhance the perception robustness of the LLM, we add Gaussian noise to the encoder’s output features to simulate the semantic discrepancies between different modalities, inspired by Chen et al. (2024b). After two stages of pure text training, our GreenPLM acquires the ability to comprehend raw 3D point clouds by directly replacing the text encoder $f_{text}$ with a point encoder $f_{pc}$ without weight tuning.
@@
Trainable Layers & Data: Similar to Stage II, the weights of projector $f_{proj}$ here are copied from the previous stage and then still kept trainable. We continue using LoRA (Hu et al. 2021) to train $f_{LLM}$ for efficient point-LLM alignment. The point cloud encoder $f_{pc}$ remain frozen. In this stage, we train using only a small amount of 3D-text pairs (90K).
@@
$\bm{T}_{m}$ $=\text{MaxPool}(\bm{T}_{p}),$ $\quad\bm{T}_{pc}^{p}$ $=\text{SoftMax}({\bm{T}_{p}\bm{T}_{m}^{T}})\bm{T}_{p},$
where $\bm{T}_{pc}\in\mathbb{R}^{N\times 1\times C}$ is the output point token sequence of the point cloud encoder ($N=512$), $\bm{T}_{c}\in\mathbb{R}^{M\times 1\times C}$ is the central token gained via farthest point sampling (FPS) from $T_{pc}$ ($M=32$), and $\bm{T}_{p}\in\mathbb{R}^{M\times K\times C}$ represents the K-Nearest Neighborhood (KNN) tokens of $\bm{T}_{c}$ within $\bm{T}_{pc}$ ($K=8$). Then, we pass $\bm{T}_{p}$ to Max Pooling on the $K$ dimension to get $\bm{T}_{m}\in\mathbb{R}^{M\times 1\times C}$. Finally, we use cross-attention in Equ.(6) to aggregate information from $\bm{T}_{pc}\in\mathbb{R}^{512\times 1\times C}\to\bm{T}_{pc}^{p}\in\mathbb{R}^{32\times 1\times C}$. Finally, we obtain the compressed token $\bm{T}_{pc}^{p}$ using zero trainable parameters. Notely, the $\bm{T}_{pc}$ input to 0M-Pooling is from the point encoder’s second-to-last layer.
@@
An efficient and accurate model evaluation method is a shared goal in the MLLM community. We observe that existing evaluation approaches often rely on GPT-4 and GPT-3.5 to assess the similarity between generated results and ground truth sentences. While this method provides accurate evaluations, it has two major drawbacks: inconsistent API versions and high evaluation costs. For instance, the GPT-3.5-turbo-0613 model used in PointLLM and MiniGPT-3D is no longer maintained, making it difficult to replicate the results. To address these issues, we propose a new benchmark based on open-source models and introduce a new metric to evaluate data efficiency. Specifically, we use two prompts for the classification task: an Instruction-type (I) prompt, “What is this?”, and a Completion-type (C) prompt, “This is an object of.”. For the captioning task, we use a single prompt: “Caption this 3D model in detail.”. We then replace GPT-4 and GPT-3.5 with the open-source Qwen2-72B-Instruct (Yang et al. 2024a) (Qwen2 for short) to evaluate the model’s output. We introduce the Accuracy-to-3D-Data Ratio (A3DR) metric to assess a model’s efficiency in utilizing 3D data, defined as follows:
@@
where Size is the size of 3D data (K), Acc is the accuracy, $\epsilon=1e-5$ to avoid zero division.
@@
InstructBLIP-7B (Dai et al. 2024) NIPS23 7B 0K Single-V. Img. 17.67 22.81 21.50 26.00 22.00 1.000
InstructBLIP-13B (Dai et al. 2024) NIPS23 13B 0K Single-V. Img. 21.56 21.92 21.50 21.50 21.62 1.000
LLaVA-1.5-7B (Liu et al. 2024) CVPR24 7B 0K Single-V. Img. 27.11 21.68 37.50 30.00 29.07 1.000
LLaVA-1.5-13B (Liu et al. 2024) CVPR24 13B 0K Single-V. Img. 27.71 27.76 39.50 35.50 32.62 1.000
GPT-4o mini (Jacob et al. 2024) OpenAI - 0K Single-V. Img. 22.00 23.10 39.00 35.00 29.78 1.000
@@
PointLLM-7B (Xu et al. 2023) ECCV24 7B 90K Point Cloud 45.22 39.30 59.00 53.00 49.13 0.749
MiniGPT-3D (Tang et al. 2024) MM24 2.7B 90K Point Cloud 43.56 43.03 54.50 55.00 49.02 0.748
GreenPLM (Ours) - 3.8B 90K Point Cloud 58.95 (+13.73) 62.36 (+19.33) 60.50 (+1.50) 58.50 (+3.50) 60.08 (+10.95) 0.792
@@
GPT4Point (Qi et al. 2024b) CVPR24 2.7B 660K Point Cloud 21.39 21.07 49.00 46.50 34.49 0.526
PointLLM-7B (Xu et al. 2023) ECCV24 7B 730K Point Cloud 51.34 50.36 62.00 63.00 56.68 0.539
PointLLM-13B (Xu et al. 2023) ECCV24 13B 730K Point Cloud 51.70 52.67 61.50 63.00 57.22 0.539
MiniGPT-3D (Tang et al. 2024) MM24 2.7B 730K Point Cloud 61.99 60.49 65.00 68.50 64.00 0.544
Table 3: 3D object captioning results on Objaverse. The results are from Qwen2 evaluation, and traditional metrics.
@@
InstructBLIP-7B (Dai et al. 2024) NIPS23 7B 0K Single-V. Img. 16.10 35.79 36.67
InstructBLIP-13B (Dai et al. 2024) NIPS23 13B 0K Single-V. Img. 13.79 33.52 35.60
LLaVA-1.5-7B (Liu et al. 2024) CVPR24 7B 0K Single-V. Img. 17.80 39.32 41.08
LLaVA-1.5-13B (Liu et al. 2024) CVPR24 13B 0K Single-V. Img. 16.00 39.64 40.90
GPT-4o mini (Jacob et al. 2024) OpenAI - 0K Single-V. Img. 26.00 38.70 39.13
@@
As shown in Tab. 2, our GreenPLM-0 achieves an average classification accuracy (AvgAcc) of 54.57% without using any 3D data, outperforming all 2D-based models. It surpasses LLaVA-1.5-13B by +21.95 and Point-Bind LLM by +27.89 in AvgAcc. Remarkably, our model also exceeds GPT4Point (660K), which is trained with 660K 3D data, by +20.08 and performs on par with PointLLM-7B (730K). With only a small amount of 3D data (90K), GreenPLM achieves an average accuracy of 60.08%, surpassing PointLLM and MiniGPT by +10.95 and +11.06 in AvgAcc, respectively. GreenPLM even outperforms PointLLM-13B (730K) while using a smaller LLM, and obtains results comparable to SoTA model MiniGPT-3D (730K). Additionally, GreenPLM (90K) outperforms MiniGPT-3D (90K) and MiniGPT-3D (730K) on the A3DR (average accuracy) by +5.9% and 45.6%, respectively. These results demonstrate the high 3D data-efficiency of our model.
3D Object Captioning
Table 4: Ablation on 3-Stage Training and Token Fusion.
@@
Sample [Uncaptioned image][Uncaptioned image]
Prompt What is this?
Label Guitar
Ins-BLIP telescope
P-Bind This is a bird flying in the sky.
@@
Fig. 1 and Tab. 5 present the qualitative results. As shown in Fig. 1, whether trained on text-only or with minimal 3D data, GreenPLM provides accurate, context-aware responses in multi-turn conversations. Tab. 5 shows that our GreenPLM-0 effectively identifies objects and understands details like color and components with text-only data. 2D-based methods like Instruct-BLIP (Ins-BLIP) and GPT-4o mini lose 3D information, suffering from occlusion, ambiguity and severe hallucinations. Point-Bind LLM (P-B LLM) lacks accurate 3D perception due to its non-robust 3D-2D-LLM connection. While using few 3D data (90K), GreenPLM offers significantly more detailed descriptions and better captures local details in point clouds, such as guitar strings and octopus suction cups, compared to PointLLM.
@@
Figure 6: Ablation on 0M-Pooling.
Figure 7: Ablation on T3D caption.
@@
Figure 8: Ablation on Text data size in Stage I & II.
Figure 9: Daily text data.
Figure 10: Noise Std.

### v3

@@ Abstract
Enabling Large Language Models (LLMs) to comprehend the 3D physical world remains a significant challenge. Due to the lack of large-scale 3D-text pair datasets, the success of LLMs has yet to be replicated in 3D understanding. In this paper, we rethink this issue and propose a new task: 3D Data-Efficient Point-Language Understanding. The goal is to enable LLMs to achieve robust 3D object understanding with minimal 3D point cloud and text data pairs. To address this task, we introduce GreenPLM, which leverages more text data to compensate for the lack of 3D data. First, inspired by using CLIP to align images and text, we utilize a pre-trained point cloud-text encoder to map the 3D point cloud space to the text space. This mapping leaves us to seamlessly connect the text space with LLMs. Once the point-text-LLM connection is established, we further enhance text-LLM alignment by expanding the intermediate text space, thereby reducing the reliance on 3D point cloud data. Specifically, we generate 6M free-text descriptions of 3D objects, and design a three-stage training strategy to help LLMs better explore the intrinsic connections between different modalities. To achieve efficient modality alignment, we design a zero-parameter cross-attention module for token pooling. Extensive experimental results show that GreenPLM requires only 12% of the 3D training data used by existing state-of-the-art models to achieve superior 3D understanding. Remarkably, GreenPLM also achieves competitive performance using text-only data.
Codehttps://github.com/TangYuan96/GreenPLM
Introduction
@@
To address this data-limited multimodal alignment problem, we propose GreenPLM. Intuitively, as shown in Fig. 2, we observe that after establishing the point-text-LLM connection, instead of increasing point-text data pairs to optimize the point-text mapping like in existing methods (Xu et al. 2023; Qi et al. 2024a), we can also enhance the text-LLM alignment by simply adding more text data. This approach can also improve the point-LLM alignment and, more importantly, reduce the reliance on point-text data pairs, shifting the data bottleneck from expensive and scarce 3D-text data to abundant and cheap text data. That is, the text-LLM alignment method fits perfectly with the goal of 3D data-efficient point-language understanding, also offers an alternative solution for aligning point clouds with LLMs, enabling GreenPLM to achieve robust 3D understanding even with limited 3D data.
@@
Caption 1M A 3D model of a friendly white dog…
@@
The encoder maps raw data into a more compact embedding space, which can then be aligned with LLMs. To reduce the training cost, one can intuitively employ a multimodal pre-trained encoder, such as CLIP (Radford et al. 2021), which has been trained on text-image pairs, for aligning 2D images with LLMs. This makes it easier to align data from different modalities with LLMs. Similarly, some existing 3D-LLMs use multimodal pre-trained encoders (Huang et al. 2023; Xue et al. 2023; Qi et al. 2023; Gao et al. 2024; Chen et al. 2024a) to map point clouds into embedding space, followed by fine-tuning the 3D-LLM. However, even without training the encoder, constructing the 3D-LLM still requires a vast amount of point-text data (Xu et al. 2023; Zhou et al. 2023; Qi et al. 2024a; Tang et al. 2024). We observe that existing methods underutilize the potential of the text encoder, only focusing on aligning point encoder with LLM. In contrast, we propose leveraging the cost-efficient text space and the text encoder to reduce the dependency on 3D data.
@@
Notably, to enhance the perception robustness of the LLM, we add Gaussian noise to the encoder’s output features to simulate the semantic discrepancies between different modalities, inspired by Chen et al. (2024b). After two stages of pure text training, our GreenPLM acquires the ability to comprehend raw 3D point clouds by directly replacing the text encoder $f_{text}$ with a paired point encoder $f_{pc}$ from Uni3D (Zhou et al. 2023) without weight tuning.
@@
Trainable Layers & Data: Similar to Stage II, the weights of projector $f_{proj}$ here are copied from the previous stage and then still kept trainable. We continue using LoRA (Hu et al. 2021) to train $f_{LLM}$ for efficient point-LLM alignment. The normalization layers and MLP in Uni3D (Zhou et al. 2023) used to align the point and text encoders are fine-tuned for 0M-Pooling output, with other weights frozen. In this stage, we train using only a small amount of 3D-text pairs (90K).
@@
$\bm{T}_{m}$ $=\text{MaxPool}(\bm{T}_{p}),$ $\quad\bm{T}_{pc}^{p}$ $=\text{SoftMax}(({\bm{T}_{p}\bm{T}_{m}^{T}})^{T})\bm{T}_{p},$
where $\bm{T}_{pc}\in\mathbb{R}^{N\times C}$ ($N=512$) is the output point token sequence of the point cloud encoder, $\bm{T}_{c}\in\mathbb{R}^{M\times C}$ ($M=32$) is the central token gained via farthest point sampling (FPS) from $T_{pc}$, and $\bm{T}_{p}\in\mathbb{R}^{M\times K\times C}$ ($K=8$) represents the K-Nearest Neighborhood (KNN) tokens of $\bm{T}_{c}$ within $\bm{T}_{pc}$. Then, we pass $\bm{T}_{p}$ to Max Pooling on the $K$ dimension to get $\bm{T}_{m}\in\mathbb{R}^{M\times 1\times C}$. Finally, we use cross-attention in Equ.(6) to aggregate information from $\bm{T}_{pc}\in\mathbb{R}^{512\times C}\to\bm{T}_{pc}^{p}\in\mathbb{R}^{32\times C}$. We obtain the compressed token $\bm{T}_{pc}^{p}$ using zero trainable parameters. Notely, the $\bm{T}_{pc}$ input to 0M-Pooling is from the point encoder’s second-to-last layer.
@@
An efficient and accurate model evaluation method is a shared goal in the MLLM community. We observe that existing evaluation approaches often rely on GPT-4 and GPT-3.5 to assess the similarity between generated results and ground truth sentences. While this method provides accurate evaluations, it has two major drawbacks: inconsistent API versions and high evaluation costs. For instance, the GPT-3.5-turbo-0613 model used in PointLLM and MiniGPT-3D is no longer maintained, making it difficult to replicate the results. To address these issues, we propose a new benchmark based on open-source models and introduce a new metric to evaluate data efficiency. Specifically, we use two prompts for the classification task: an Instruction-type (I) prompt, “What is this?”, and a Completion-type (C) prompt, “This is an object of.”. For the captioning task, we use a single prompt: “Caption this 3D model in detail.”. We then replace GPT-4 and GPT-3.5 with the open-source Qwen2-72B-Instruct (Yang et al. 2024a) (Qwen2 for short) to evaluate the model’s output. We also introduce the Accuracy-to-3D-Data Ratio (A3DR) metric to assess a model’s efficiency in utilizing 3D data, defined as follows:
@@
where Size is the size of 3D data (K), Acc is the accuracy, $\epsilon=1e-5$ prevents zero division, $\lambda=3$ adjusts discrimination.
@@
InstructBLIP-7B (Dai et al. 2024) NIPS23 7B 0K Single-Img. 17.67 22.81 21.50 26.00 22.00 1.000
InstructBLIP-13B (Dai et al. 2024) NIPS23 13B 0K Single-Img. 21.56 21.92 21.50 21.50 21.62 1.000
LLaVA-1.5-7B (Liu et al. 2024) CVPR24 7B 0K Single-Img. 27.11 21.68 37.50 30.00 29.07 1.000
LLaVA-1.5-13B (Liu et al. 2024) CVPR24 13B 0K Single-Img. 27.71 27.76 39.50 35.50 32.62 1.000
GPT-4o mini (Jacob et al. 2024) OpenAI - 0K Single-Img. 22.00 23.10 39.00 35.00 29.78 1.000
@@
PointLLM-7B (Xu et al. 2023) ECCV24 7B 90K Point Cloud 45.22 39.30 59.00 53.00 49.13 0.674
MiniGPT-3D (Tang et al. 2024) MM24 2.7B 90K Point Cloud 43.56 43.03 54.50 55.00 49.02 0.673
GreenPLM (Ours) - 3.8B 90K Point Cloud 58.95 (+13.73) 62.36 (+19.33) 60.50 (+1.50) 58.50 (+3.50) 60.08 (+10.95) 0.762
@@
GPT4Point (Qi et al. 2024b) CVPR24 2.7B 660K Point Cloud 21.39 21.07 49.00 46.50 34.49 0.078
PointLLM-7B (Xu et al. 2023) ECCV24 7B 730K Point Cloud 51.34 50.36 62.00 63.00 56.68 0.116
PointLLM-13B (Xu et al. 2023) ECCV24 13B 730K Point Cloud 51.70 52.67 61.50 63.00 57.22 0.117
MiniGPT-3D (Tang et al. 2024) MM24 2.7B 730K Point Cloud 61.99 60.49 65.00 68.50 64.00 0.131
@@
InstructBLIP-7B (Dai et al. 2024) NIPS23 7B 0K Single-Img. 16.10 35.79 36.67
InstructBLIP-13B (Dai et al. 2024) NIPS23 13B 0K Single-Img. 13.79 33.52 35.60
LLaVA-1.5-7B (Liu et al. 2024) CVPR24 7B 0K Single-Img. 17.80 39.32 41.08
LLaVA-1.5-13B (Liu et al. 2024) CVPR24 13B 0K Single-Img. 16.00 39.64 40.90
GPT-4o mini (Jacob et al. 2024) OpenAI - 0K Single-Img. 26.00 38.70 39.13
@@
Table 3: 3D object captioning results on Objaverse. The results are from Qwen2 evaluation and traditional metrics.
@@
As shown in Tab. 2, our GreenPLM-0 achieves an average classification accuracy (AvgAcc) of 54.57% without using any 3D data, outperforming all 2D-based models. It surpasses LLaVA-1.5-13B by +21.95 and Point-Bind LLM by +27.89 in AvgAcc. Remarkably, our model also exceeds GPT4Point (660K), which is trained with 660K 3D data, by +20.08 and performs on par with PointLLM-7B (730K). With only a small amount of 3D data (90K), GreenPLM achieves an average accuracy of 60.08%, surpassing PointLLM and MiniGPT-3D by +10.95 and +11.06 in AvgAcc, respectively. GreenPLM even outperforms PointLLM-13B (730K) while using a smaller LLM, and obtains results comparable to SoTA model MiniGPT-3D (730K). Additionally, GreenPLM (90K) outperforms MiniGPT-3D (90K) and MiniGPT-3D (730K) on the A3DR (average accuracy) by +8.9% and 63.1%, respectively. These results demonstrate the high 3D data-efficiency of our model.
Sample [Uncaptioned image][Uncaptioned image]
Prompt What is this?
Label Guitar
Ins-BLIP telescope
P-Bind This is a bird flying in the sky.
@@
Table 5: Ablation on 3-Stage Training and Token Fusion.
3D Object Captioning
Figure 6: Ablation on 0M-Pooling.
Figure 7: Ablation on T3D caption.
Figure 8: Ablation on Text data size in Stage I & II.
Figure 9: Daily text data.
Figure 10: Noise Std.
@@
Fig. 1 and Tab. 4 present the qualitative results. As shown in Fig. 1, whether trained on text-only or with minimal 3D data, GreenPLM provides accurate, context-aware responses in multi-turn conversations. Tab. 4 shows that our GreenPLM-0 effectively identifies objects and understands details like color and components with text-only data. 2D-based methods like Instruct-BLIP (Ins-BLIP) and GPT-4o mini lose 3D information, suffering from occlusion, ambiguity and severe hallucinations. Point-Bind LLM (P-B LLM) lacks accurate 3D perception due to its non-robust 3D-2D-LLM connection. While using few 3D data (90K), GreenPLM offers significantly more detailed descriptions and better captures local details in point clouds, such as guitar strings, compared to PointLLM.
@@ Conclusion
Acknowledgments
This work was supported by the China National Natural Science Foundation No. 62202182, No. 62176101, No. 62276109, and also supported by Guangdong Basic and Applied Basic Research Foundation 2024A1515010224, 2024A1515030017 and 2024A1515011153.